# arXiv:2603.01290v1[cs.AI] 01 Mar 2026

### v1

@@ Appendix C DQN Hyperparameters
$\epsilon$ (start) 1.0
$\epsilon$ (end) 0.05
$\epsilon$ decay steps 20,000
@@ Appendix D FastF1 Observable Derivation Pipeline
The within-sector speed trace is linearly detrended by fitting and subtracting a first-order polynomial. The variance of the detrended residual is computed as $\sigma^{2}_{\text{speed}}$.
@@
In designated activation zones, Active Aero deployment is inferred from the speed residual after removing the ERS-explained component. A positive residual above 2.0 km/h is classified as $z_{\text{aero}}=1$.
@@ E.1 Strategy Profiles
Profile $p_{\text{burn}}$ schedule Trap execution Notes
@@
Trap-setter 0.30 baseline; 0.90 post-trap In activation zones when gap <1.5<1.5 s Primary test of Proposition 1
@@
All four profiles use the physics-derived transition matrices ($T^{\text{burn}}$, $T^{\text{harvest}}$) with $p_{\text{burn}}$ as the mixing weight. Tyre degradation evolves independently of strategy profile, following the monotone absorbing chain with stint-appropriate starting states. 5 races are generated per profile (20 total), each of 58 laps $\times$ 3 sectors $\times$ 6 rival cars = 1,044 sector observations per race, for 20,880 total observations across the evaluation set.
@@ E.2 Observation Noise Model
Synthetic observations are generated from the emission parameters (Section 4) with i.i.d. Gaussian noise added to each continuous observable:
$\hat{o}_{t}=o_{t}+\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(0,\sigma^{2}_{\text{noise}})$ (17)
where $\sigma_{\text{noise}}=0.2$ km/h for $\Delta v_{\text{trap}}$, $\sigma_{\text{noise}}=0.05$ s for $\Delta t_{\text{sector}}$, and $\sigma_{\text{noise}}=1$ m for $\Delta b_{\text{brake}}$. These values represent sensor noise and measurement uncertainty, set to approximately one-third of the emission standard deviations to ensure the signal-to-noise ratio reflects the real-telemetry regime. $z_{\text{aero}}$ is generated directly from the Bernoulli emission parameters with no additional noise. $\sigma^{2}_{\text{speed}}$ is generated from a truncated Gaussian to ensure non-negativity, with the truncation at 0.

### v2

@@ Appendix C DQN Hyperparameters
$\varepsilon$ (start) 1.0
$\varepsilon$ (end) 0.05
$\varepsilon$ decay steps 20,000
@@ Appendix D FastF1 Observable Derivation Pipeline
The within-sector speed trace is linearly detrended by fitting and subtracting a first-order polynomial. The variance of the detrended residual is computed as $\sigma^{2}_{\mathrm{speed}}$.
@@
In designated activation zones, Active Aero deployment is inferred from the speed residual after removing the ers-explained component. A positive residual above 2.0 km/h is classified as $z_{\mathrm{aero}}=1$.
Super-clipping duration ($\delta_{\mathrm{throttle}}$).
For each sector, the fraction of straight-line telemetry samples in which throttle $\geq 98\%$ and speed is simultaneously below the driver’s 5-lap rolling speed baseline is computed from the FastF1 car_data Throttle channel at 10 Hz. If the Throttle channel is absent, $\delta_{\mathrm{throttle}}$ defaults to the circuit-appropriate prior mean for the inferred ERS state.
@@ E.1 Strategy Profiles
Profile $p_{\mathrm{burn}}$ schedule Trap execution Notes
@@
Trap-setter 0.30 baseline; 0.90 post-trap In activation zones when gap <1.5<1.5 s Primary test of Proposition 1; operates from $L_{\mathrm{harvest}}$
@@ E.2 Observation Noise Model
Synthetic observations are generated from the emission parameters (Section 4) with i.i.d. Gaussian noise:
$\hat{o}_{t}=o_{t}+\varepsilon_{t},\quad\varepsilon_{t}\sim\mathcal{N}(0,\sigma^{2}_{\mathrm{noise}})$ (18)
where $\sigma_{\mathrm{noise}}=0.2$ km/h for $\Delta v_{\mathrm{trap}}$, $0.05$ s for $\Delta t_{\mathrm{sector}}$, and 1 m for $\Delta b_{\mathrm{brake}}$. $z_{\mathrm{aero}}$ is generated directly from Bernoulli emission parameters. $\sigma^{2}_{\mathrm{speed}}$ is generated from a truncated Gaussian (truncation at 0). $\delta_{\mathrm{throttle}}$ is generated from a truncated Gaussian on $[0,1]$ with the parameters from Table 2, with a Melbourne circuit offset (+0.10) applied to the $L_{\mathrm{derate}}$ state to reflect mandatory super-clipping.
Five races are generated per profile (20 total), each of 58 laps $\times$ 3 sectors $\times$ 6 rival cars = 1,044 sector observations per race, for 20,880 total observations.